OpenStudy (anonymous):

determine integral: (2/x-4e ^x) dx

OpenStudy (anonymous):

Integral of 2/x + Integral of -4 e^x = 2 * Integral of 1/x -4 Integral of e^x = 2* ln x - 4 e^x + C